OpenStudy (anonymous):

What's the easiest and quickest way to answer this question? Is 329341 divisible by 9?

OpenStudy (rosedewittbukater):

Divide 329341 by 9 and see if the answer is a whole number. You can use long division or a calculator.

OpenStudy (anonymous):

What if a calculator isn't accessible?

OpenStudy (rosedewittbukater):

Then use long division. Do you know how?

OpenStudy (anonymous):

You can add up the digits to see if it is divisible by 3. The test wouldn't be conclusive though.

OpenStudy (anonymous):

3+2+9+3+4+1 = 22, so in this case the test will tell you it isn't divisible by 9.

OpenStudy (anonymous):

but if they added up to 21 or something, then you would be out of luck.

OpenStudy (rosedewittbukater):

There's probably a calculator on your computer or you could find one online, or on your phone. If you mean that you aren't allowed to use a calculator, then use regular division.

OpenStudy (anonymous):

@rosedewittbukater I think the point of the question is not to use a calculator. But if I wanted to I probably would :)

OpenStudy (anonymous):

@wio @rosedewittbukater THANK YOU!

OpenStudy (rosedewittbukater):

Oh haha ok. No problem!

OpenStudy (anonymous):

only divisible by 9 if the digits add up to 9 ex: 18= 1+8=9 72=7+2+9

OpenStudy (anonymous):

@Luis_Rivera the whole set of rules for the following numbers: 2, 3, 4, 5, 9, or 10?

OpenStudy (anonymous):

if numer even- divisible by 2 and 4 if digits ad up to 3 - divisible by 3 if number ends in 5- divisible only by 5 if number ends in 0 ir 5- divisible by 10 and 5

OpenStudy (anonymous):

@Luis_Rivera Thanks

OpenStudy (anonymous):

kk